List of publications of András RECSKI

February 20, 2008

1. Recski A.: A Fibonacci-számok általánosításáról, Matematikai Lapok 22 (1971) 165-170.

2. L. Lovász and A. Recski: On the sum of matroids, Acta Math. Acad. Sci. Hungary 24 (1972) 329-333.

3. Recski A.: Matroidok alkalmazása a villamos hálózatok analízisében, Matematikai Lapok 24 (1973) 311-328.

4. A. Recski: On partitional matroids with applications, Coll. Math. Soc. J. Bolyai 10 (Infinite and Finite Sets, North-Holland, Amsterdam, 1974) III. 1169-1179.

5. A. Recski: Enumerating partitional matroids, Studia Sci. Math. Hungar. 9 (1974) 247-249.

6. Á. Csurgay, Z. Kovács and A. Recski: Transient analysis of lumped-distributed non-linear networks, Proc. 5th Internat. Coll. on Microwave Communication, Budapest, 1974.

7. A. Recski: On the generalization of the Fibonacci numbers, The Fibonacci Quarterly 13 (1975) 315-317.

8. Recski A.: Matroidok alkalmazása a villamos hálózatok analízisében II. Matematikai Lapok 26 (1975) 287-303.

9. A. Recski: Matroids and state variables, Proc. 1976 European Conf. on Circuit Theory and Design, Genova, 1976, I. 44-51.

10. A. Recski: On the sum of matroids II, Proc. 5th British Combinatorial Conf. Aberdeen (Utilias Math., Winnipeg, 1976) 515-520.

11. A. Recski: On random partitions, Discrete Math. 16 (1976) 173-177.

12. A. Recski: Application of graph theory to network analysis – a survey, Beiträge zur Graphentheorie und Anwendungen, Oberhof (GDR), 1977, 193-200.

13. A. Recski: Matroids and independent state variables, Annual of the Research Institute for Telecommunication, Budapest, 1977, II. 239-246.

14. A. Recski: Sufficient conditions for the unique solvability of networks containing linear memoryless 2-ports, Proc. 1978 European Conf. on Circuit Theory and Design, Lausanne, 1978, I. 93-97.

15. A. Recski: Contributions to the n-port interconnection problem by means of matroids, Coll.

Math. Soc. J. Bolyai 18 (Combinatorics, North-Holland, Amsterdam, 1978) II. 877-892.

16. A. Recski: Matroids in network theory, Proc. 6th Internat. Coll. on Microwave Communication, Budapest, 1978, I.II-2/12.1-4.

17. A. Recski: Decompositions of a graphic matroid, Mathématiques discrètes: Codes et hypergraphes (Bruxelles), Cahiers du CERO, 20 (1978) 437-442.

18. A. Recski: Matroidal structure of n-ports, Coll. Math. Soc. J. Bolyai 18 (Combinatorics, North-Holland, Amsterdam, 1978) II. 893-909.

19. A. Recski: Unique solvalibity and order of complexity of linear networks containing memoryless n-ports, Circuit Theory and Applications 7 (1979) 31-42.

20. A. Recski: Terminal solvability and the nport interconnection problem, Proc. IEEE Internat. Symp. Circuits and Systems, Tokyo, 1979, 988-991. 21.

Proc. IECEJ Coll. Cir-

cuits and Systems, CAS-79-78 (1979) 19-25.

22. M. Iri and A. Recski: Reflections on the concepts of dual, inverse and adjoint networks, Research memorandum RMI 97-7, University of Tokyo, September, 1979.

23. A. Recski: Some remarks on the arboricity of the tree complements, Trans. Fac. Sci. Tokai University 15 (1979) 71-74.

24. A. Recski: Sufficient conditions for the unique solvability of networks containing linear memoryless 2-ports, Circuit Theory and Applications 8 (1980) 95-103.

25. M. Iri and A. Recski: What does duality really mean? Circuit Theory and Applications 8 (1980) 317-324.

26. A. Recski: Matroids and network synthesis, Proc. 1980. European Conf. on Circuit Theory and Design, Warsaw, 1980, II. 192-197.

27. A. Recski and M. Iri: Network theory and transversal matroids, Discrete Applied Math. 2 (1980) 311-326.

Proc. IECEJ Coll. Circuits and Systems, CAS-79-133 (1980) 5-10.

29. M. Iri and A. Recski: Reflections on the concepts of dual, inverse and adjoint networks II, Towards a qualitative theory, Research memorandum RMI 80-1, University of Tokyo, January 1980.

30. A. Recski: An algorithm to determine whether the sum of some graphic matroids is graphic, Coll. Math. Soc. J. Bolyai 25 (Algebraic methods in graph theory, North-Holland, Amsterdam, 1981) II. 647-656.

31. A. Recski and V. Zoller: On the parametrization of linear memoryless 2-ports, Proc. 1981 European Conf. on Circuit Theory and Design, The Hague, 1981, 758-763.

32. A. Recski and J. Takács: On the combinatorial sufficient conditions for linear network solvability, Circuit Theory and Applications 9 (1981) 351-354.

33. A. Recski: On the sum of matroids III, Discrete Math. 36 (1981) 273-287.

34. A. Recski: A practical remark on the minimal synthesis of resistive n-ports, IEEE Trans. Circuits and Systems CAS-29 (1982) 267-269.

35. M. Iri and A. Recski: Duality and reciprocity – a qualitative approach, Proc. IEEE Internat. Symp. Circuits and Systems, Rome, 1982, II, 415-418.

36. A. Recski: Matroids and terminal solvability, Proc. 7th Internat. Coll. on Microwave Communication, Budapest, 1982, I. 242-245.

37. A. Recski and V. Zoller: On the parametrization of linear memoryless 2-ports, Circuit Theory and Application 10 (1982) 57-67.

38. A. Recski: Engineering applications of matroids – a survey, A. Barlotti (ed): Matroid Theory and Its Applications, Liguori editore, Napoli, 1982, 299-321.

39. L. Lovász and A. Recski: Selected topics of matroid theory and its applications, Rendiconti del Circolo Matematico di Palermo II. 2 (1982) 171-185.

40. A. Recski: On the generalization of the matroid parity and the matroid partition problems, with applications, Annals of Discrete Math. 17 (1983) 567-574.

41. A. Recski: Unique solvability of linear passive memoryless networks – a survey, Proc. 1983 European Conf. on Circuit Theory and Design, Stuttgart, 1983, 53-55.

42. A. Recski: Local and global inconsistencies in the n-port interconnection problem, Circuit Theory and Applications 11 (1983) 317-375.

43. A. Recski: Some recent results on planarity and duality, Proc. of the graph theory conference, dedicated to the memory of K. Kuratowski, Łagow, Poland, 1981, Springer Lecture Notes 1018 (1983) 199-213.

44. A. Recski: A network theory approach to the rigidity of skeletal structures I. Modelling and interconnection, Discrete Applied Math. 7 (1984) 313-324.

45. A. Recski: Statics and electric network theory: A unifying role of matroids, W. R. Pulleyblank (ed): Progress in Combinatorial Optimization, Academic Press, London, 1984, 307-314.

46. A. Recski: A network theory approach to the rigidity of skeletal structures II. Laman’s theorem and topological formulæ, Discrete Applied Math. 8 (1984) 63-68.

47. A. Recski: Applications of combinatorics to statics – a survey, Rendiconti del Circolo Matematico di Palermo II. 3 (1984) 237-247.

48. A. Recski: A network theory approach to the rigidity of skeletal structures III. An electric model for planar frameworks, Structural Topology 9 (1984) 59-71.

49. A. Recski: Unique solvability of linear memoryless networks – a survey, IEEE Trans. Circuits and Systems CAS-31 (1984) 894-897.

50. A. Recski: Some problems of self-dual matroids, Coll. Math. Soc. J. Bolyai 37 (Finite and Infinite Sets, North-Holland, Amsterdam, 1984) 635-648.

51. L. Lovász and A. Recski (eds): Matroid Theory (Proceedings of the matroid theory colloquium in Szeged, Hungary, 1982) North-Holland, Amsterdam, 1985.

52. A. Recski: Some open problems of matroid theory, suggested by its applications, Coll. Math. Soc. J. Bolyai 40 (Matroid theory, North-Holland, Amsterdam, 1985) 311-325.

53. N. Chakravarty, G. Holman, S. McGuinness and A. Recski: One-story buildings as tensegrity frameworks, Structural Topology 12 (1986) 11-18.

54. A. Recski: Interconnection properties and matroidal characterization of some basic concepts of network theory, Proc. 8th Internat. Coll. on Microwave Communication, Budapest, 1986, 47.

55. A. Recski: Elementary strong maps of graphic matroids, Graphs and Combinatorics 3 (1987) 379-382.

56. A. Recski: Engineering applications of matroids – a survey, Proc. Conference on Control Systems and Computer Science, Bucharest, 1987, 180-183.

57. A. Recski: Some algorithmic problems of linear network solvability, Proc. IEEE Internat. Symp. Circuits and Systems, Helsinki, 1988, 135-138.

58. A. Recski: Bracing cubic grids – a necessary condition, Discrete Mathematics 73 (1988/89) 199-206.

59. A. Recski: Some open problems in matroid theory (Combinatorial Mathematics: Proc. 3rd Internat. Conference), Annals of the New York Academy of Sciences 555 (1989) 332-334.

60. A. Recski: Matroid theory and its applications in electric network theory and in statics (Algorithms and Combinatorics, Vol. 6) Springer Verlag, Berlin (ISBN 3-540-15285-7), New York (ISBN 0-387-15285-7) and Akadémiai Kiadó (ISBN 963-05-5253-1), Budapest, 1989.

61. A. Recski: Symmetric bracing of one-story buildings with cables and asymmetric bracing of one-story buildings with rods, Symmetry of structures, Budapest, 1989, Vol. II. 471-472.

62. A. Recski and F. Stryzewski: Vertex-disjoint channel routing on two layers, in Ravi Kannan and W. R. Pulleyblank (eds.): Integer Programming and Combinatorial Optimization, University of Waterloo Press, Waterloo, Ont., 1990, 397-405.

63. A. Recski: Változások a BME Villamosmérnöki Karán a Matematika tárgy oktatásában, Híradástechnika 42 (1991. június) 35.

64. A. Recski: One-story buildings as tensegrity frameworks II. Structural Topology 17 (1991) 43-52.

65. A. Recski: Applications of Combinatorics in Structural Engineering, First Workshop on Applications of Combinatorial Optimization in Science and Technology, Rutgers University, New Brunswick, NJ. 1991, 275-282.

66. B. Andrásfai, P. Ablonczy, A. Recski and K. Vesztergombi: Teaching Discrete Mathematics to Undergraduate Engineering Students, Proc. of the 6th European Seminar on Mathematics in Engineering Education, Budapest–Balatonfüred, Hungary, 1991, 6-10.

67. A. Recski: Minimax results and polynomial algorithms in VLSI routing, in M. Fiedler and J. Nešetřil, eds., Combinatorics, Graphs, Complexity (Proc. 4th Czechoslovak Symposium on Combinatorics, Prachatice, 1990), Annals of Discrete Math. (1992) 261-273.

68. A. Recski: Applications of combinatorics to statics – a second survey, in J. Nešetřil (ed): Topological, Algebraical and Combinatorial Structures, Discrete Math. 108 (1992) 183-188.

69. A. Recski: Hogyan tegyük tönkre konferencián tartott előadásainkat az írásvetítő alkalmazásával? Matematikai Lapok 2 (1992) 2, 14-16.

70. A. Recski and W. Schwärzler: One-story buildings as tensegrity frameworks III, Discrete Applied Math. 39 (1992) 137-146.

71. A. Recski: Some new developments in the undergraduate math curriculum, Proc. of the 7th European Seminar on Mathematics in Engineering Education, Eindhoven, The Netherlands, 1993, 39. 72. G. Y. Katona and A. Recski: Introduction to Finite Mathematics (University textbook in Hungarian), L. Eötvös University Press, Budapest, 1993.

73. Császár Ákosné, Recski András and Rózsa Pál: A BME Villamosmérnökkari Matematika Tanszéke, Elektronikai Technológia – Mikrotechnika, 30 (1993) 209-211.

74. G. Y. Katona and A. Recski: Combinatorics, Graph Theory and the Theory of Algorithms (University textbook in Hungarian), Technical University of Budapest, Budapest, 1993.

75. A. Recski: Elementary strong maps of graphic matroids II, Graphs and Combinatorics 10 (1994) 205-206.

76. A. Recski: Combinatorics in electrical engineering and statics, in R. Graham, M. Grötschel and L. Lovász (eds.): Handbook in Combinatorics, Elsevier, Amsterdam, 1995, pp. 1911-1924.

77. E. Boros, A. Recski and F. Wettl: Unconstrained multilayer switchbox routing, Annals of Operations Research 58 (1995) 481-491.

78. T. Jordán and A. Recski: Combinatorial Optimization (University textbook in Hungarian), Technical University of Budapest and L. Eötvös University, Budapest, 1995.

79. A. Recski: Calculus exercises with and without Mathematica, Yale University, New Haven, 1995.

80. A. Recski: Some polynomially solvable subcases of the detailed routing problem in VLSI design, Proc. Oper. Research Conf., Springer, Berlin, 1997, pp. 107-110.

81. A. Recski: Channel routing in the doglegfree multilayer Manhattan model, Proc. 1997. European Conf. on Circuit Theory and Design, Budapest, 1997, I. 39-43.

82. Zs. Gáspár, N. Radics and A. Recski: Square grids with long diagonals, Optimization Methods and Software, 10 (1998) 217-231.

83. Gy. Nagy and A. Recski: Rúd-csukló szerkezetek, Matematikai Lapok, 1998. 72-75.

84. A. Balog, G. O. H. Katona, A. Recski and D. Szász (eds.): European Congress of Mathematics, Proceedings. Birkhäuser Verlag, Basel, 1998.

85. Zs. Gáspár, N. Radics and A. Recski: How to make square grids rigid? 10th Inter-Institute Seminar on Nonlinear Computational Mechanics, (Technical University of Budapest and Cracow University of Technology), Budapest, 1998, 22.

86. L. Lovász, A. Gyárfás, Gy. Katona, A. Recski and L. Székely (eds.): Graph theory and combinatorial biology, Bolyai Society Mathematical Studies 7., Budapest, 1999.

87. E. Boros, A. Recski, T. Szkaliczki and F. Wettl: Polynomial time Manhattan routing without doglegs – a generalization of Gallai’s algorithm, Computers and Artificial Intelligence, 18 (1999) 403-413.

88. Zs. Gáspár, N. Radics and A. Recski: Rigidity of square grids with holes, Computer Assisted Mechanics and Engineering Sciences, 6 (1999) 329-335.

89. A. Recski and D. Szeszlér: 3-dimensional single active layer routing, in J. Akiyama, M. Kano and M. Urabe (eds.): Discrete and Computational Geometry, Lecture Notes in Computer Science, Springer, Vol. 2098. 2000, 318-329.

90. N. Katoh, T. Ibaraki and A. Recski (eds.): Selected Papers, First Japanese-Hungarian Symposium for Discrete Mathematics and its Applications, Discrete Applied Math. 115 (2001) pp. 1-222.

91. Z. Mann, A. Orbán and A. Recski: Aufgaben zur theoretischen Informatik (University textbook in German), Budapest University of Technology and Economics, Budapest, 2001.

92. J. Tapolcai, P. Laborczi, Pin-Han Ho, A. Recski, T. Cinkler and H. T. Mouftah: Algorithms for asymmetrically weighted pair of disjoint paths in survivable optical networks, Proc. 3rd Internat. Workshop on Design of Reliable Communication Networks, Budapest, 2001. pp. 228-235.

93. A. Recski: Some polynomially solvable subcases of the detailed routing problem in VLSI design, Discrete Applied Math. 115 (2001) 199-208.

94. Gy. Y. Katona, A. Recski and Cs. Szabó: Foundation of Computer Science (University textbook in Hungarian), Typotex, Budapest (ISBN 963-9326-24-0) 2002.

95. Recski A.: Dualitás a matematikában és sok más helyen, in Hraskó András (ed.): Új matematikai mozaik, Typotex, Budapest, 2002, 413-426.

96. A. Recski: Two matroidal families on the edge set of a graph, Discrete Math. 251 (2002) 155-162.

97. N. Radics and A. Recski: Applications of combinatorics to statics – rigidity of grids, Discrete Applied Math. 123 (2002) 473-485.

98. P. Laborczi and A. Recski: Graph theory and its applications in communication networks, Chapter 1.10 of the on-line book Telecommunication networks and information services,

www.hte.hu/onlinebook.html (2002) pp. 157-

167.

99. A. Recski, G. Salamon and D. Szeszlér: Improving size-bounds for subcases of squareshaped switchbox routing, Periodica Polytechnica, 48 (2004) 55-60.

100. B. Golda, B. Laczay, Z. Á. Mann, Cs. Megyeri, A. Recski: Implementation of VLSI Routing Algorithms, W. Elmenreich (ed.) Intelligent Systems at the Service of Mankind, Ubooks, 2004, 349-360.

101. T. Jordán, A. Recski and D. Szeszlér: System Optimization (University textbook in Hungarian), Typotex (ISBN 963-9548-39-1), Budapest, 2004.

102. K. Ambrus Somogyi and A. Recski: On the complexity of the channel routing problem in the dogleg-free multilayer Manhattan model, Acta Polytechnica Hungarica, Vol. 1. Issue 2. (2004) 89-97.

103. A. Recski and O. Shai: One-dimensional synthesis of graphs as tensegrity frameworks, 4th Japanese-Hungarian Symposium on Discrete Mathematics and its Applications, Budapest, 2005, pp. 284-288.

104. A. Recski: Maps of matroids with applications, Discrete Math. 303 (2005) 175-185.

105. B. Drága and A. Recski: A new worst-case lower bound for the width of single row routing in the unconstrained two-layer model, Proc. 6th Internat. Symp. Computational Intelligence, Budapest Tech., 2005, pp. 172-176.

106. A. Recski and D. Szeszlér: The evolution of an idea – Gallai’s algorithm, Bolyai Society Mathematical Studies 15 (2006) 317-328.

107. A. Recski and J. Szabó: On the generalization of the matroid parity problem, in J. A. Bondy, J. Fonlupt, J. L. Fouquet, J.-C. Fournier and J. Ramirez Alfonsin (eds.): Graph Theory, Trends in Mathematics, Birkhauser, 2006, pp. 347-354.

108. K. Friedl, A. Recski and G. Simonyi: Graph Theory Exercises, (University textbook in Hungarian), Typotex (ISBN 963-9664-01-4), Budapest, 2006.

109. A. Recski and D. Szeszlér: Routing vertexdisjoint Steiner trees in a cubic grid and connections to VLSI, Discrete Applied Math. 155 (2007) 44-52.

110. A. Recski: Some matroidal results inspired by electric engineering applications, 5th Japanese-Hungarian Symposium on Discrete Mathematics and its Applications, Sendai, 2007, pp. 1-10

111. A. Recski and D. Szeszlér: 3-dimensional routing, 5th Japanese-Hungarian Symposium on Discrete Mathematics and its Applications, Sendai, 2007, pp. 138-145.

112. L. Szőllősi, T. Marosits, G. Fehér and A. Recski: Fast digital signature algorithm based on subgraph isomorphism, in Feng Bao et al. (eds.): Criptology and Network Security, Lecture Notes in Computer Science, Springer, Vol. 4856. 2007, 34-46.

113. A. Recski and O. Shai: Tensegrity frameworks in the one-dimensional space, to appear.

114. A. Recski: Combinatorial conditions for the rigidity of tensegrity frameworks, to appear.

115. T. Jordán, A. Recski and Z. Szabadka: Rigid tensegrity labelling of graphs, to appear.

116. A. Recski, G. Salamon and D. Szeszlér: Improving size-bounds for subcases of square-shaped switchbox routing, to appear.

117. A. Recski: Matroids – the engineers’ revenge, to appear.